. A ball is thrown horizontally from the roof of a building 45m tall and lands 24m from the base. What was the balls initial speed?


1
Expert's answer
2021-06-18T07:54:54-0400

Let U be the initial horizontal velocity of ball.

Since the horizontal velocity of the ball remains constant.

"U=\frac{Horizontal distance}{Time}"

The time taken by the ball to fall from height of H=45 m is given by kinematical equation in vertical direction.

"H=\frac{1}{2}×g×t^2\\t=\sqrt{\frac{2×H}{g}}\\t=\sqrt{\frac{2×45}{9.8}}=3.03 sec"

Hence, U=24/3.03=7.92m/s

Answer: The initial velocity is 7.92 m/s